A man, on tour, travels first 160 km at 64 km/hr and the next 160 km at 80 km/hr. the average speed for the first 320 km of the tour, is:

Total time taken  = (160/64 + 160/8)hrs

              = 9/2 hrs.



Average speed  = (320 x 2/9) km.hr

             = 71.11 km/hr.



how do we get 9/2 can I please know it as soon as possible



1
Expert's answer
2021-05-17T17:34:25-0400
"T=t_1+t_2\\T=\frac{s_1}{v_1}+\frac{s_2}{v_2}\\T=\frac{160}{64}+\frac{160}{80}=4.5\ h"